OpenStudy (rayford):

Multiply. (x – 4)(x2 – 3x + 5)

OpenStudy (neo92):

(x-4)(x^2 - 3x + 5) =x(x^2-3x+5)-4(x^2-3x+5) =x^3 - 3x^2 + 5x -4x^2 + 12x - 20 =\[x ^{3}-7x^{2}+17x-20\]

OpenStudy (rayford):

i see what you did there :]

OpenStudy (anonymous):

\[(x-4)(x^{2} -3x +5) = x^{3} - 3x^{2} +5x -4x^{2} +12x -20\] \[= x^{3} -7x^{2} + 17x -20\]